# Arcade Processes for Informed Martingale Interpolation

### v2

@@ Abstract
Arcade processes are a class of continuous stochastic processes that interpolate in a strong sense, i.e., omega by omega, between zeros at fixed pre-specified times. Their additive randomization allows one to match any finite sequence of target random variables, indexed by the given fixed dates, on the whole probability space. The randomized arcade processes (RAPs) can thus be interpreted as a generalization of anticipative stochastic bridges. The filtrations generated by these processes are utilized to construct a class of martingales which interpolate between the given target random variables. These so-called filtered arcade martingales (FAMs) are almost-sure solutions to the martingale interpolation problem and reveal an underlying stochastic filtering structure. In the special case of conditionally-Markov randomized arcade processes, the dynamics of FAMs are informed by Bayesian updating. The same ideas are applied to filtered arcade reverse-martingales, which are constructed in a similar fashion, using reverse-filtrations of RAPs, instead. As a potential application of this theory, optimal transportation is explored: FAMs may be used to introduce noise in martingale optimal transport, in a similar fashion to how Schrödinger’s problem introduces noise in optimal transport. This information-based approach to transport is concerned with selecting an optimal martingale coupling for the target random variables under the influence of the noise that is generated by an arcade process, and suggests application in finance or climate science, for instance.
Keywords: Stochastic interpolation, stochastic bridges, martingale interpolation, martingale fitting, optimal transport, nonlinear stochastic filtering, information-based approach, Markov processes, Bayesian updating.
AMS2020: 49Q22, 60G35, 60G44, 60J60, 62M20, 94A15.
@@ 1 Introduction
Strassen’s and Kellerer’s theorems do not provide a method for constructing the matching martingales. Many techniques are dedicated to Kellerer’s setting, where one is looking to construct a martingale that matches a given peacock [11]: Skorokhod’s embedding problem, Brownian random time-changes, local volatility models, scaling peacocks method, etc., see [21]. Since the target is a dynamical measure, the matching is produced in a weak sense, and the martingale evolves in continuous time. On the other hand, the matching in Strassen’s setting can be weak or almost sure, depending on the nature of the targets (measures or random variables), and the martingale is in discrete time. A widely known approach for this task is martingale optimal transport (MOT, [2], [3]). Inspired by Kantorovich’s optimal transport theory ([13], [31], [5]), it provides a framework for selecting a martingale coupling by minimizing a cost functional over the set of all martingale couplings.
@@
Our interest lies at the intersection of Strassen’s and Kellerer’s settings: how to construct a continuous-time martingale that matches, in law or almost surely, a given finite set of convexly ordered random variables, at pre-specified dates? This is for instance of interest in a financial setting, where knowledge of the prices of vanilla call and put options provides an implied distribution for the underlying asset price at future fixed dates under a no-arbitrage requirement. This is an example of what is called the martingale interpolation problem, since one wishes to interpolate with a continuous-time martingale between the components of a discrete-time martingale. Imposing the martingale property on the interpolating process is the challenging part of this problem. If one put the martingale condition aside, then bridging the gap between a coupling and an interpolating measure is usually a manageable proposition. In Kantorovich’s optimal transport theory, for example, it is well-known that there is a one-to-one correspondence between optimal couplings and optimal interpolating measures, that is, geodesics in Wasserstein spaces, called McCann’s interpolations, under mild conditions. However, this property is not present in the martingale counterpart of optimal transport. Solutions to the martingale interpolation problem exist. The martingale Benamou-Brenier problem (MBBP) [1] aims to solve the optimization problem
$\sup\limits_{\mathcal{M}_{t}(\mu,\nu,B)}\mathbb{E}\left[\int_{0}^{1}\sigma_{t}\,\mathrm{d}t\right]$
over the set
$\mathcal{M}_{t}(\mu,\nu,B)$ =$=\left\{(\sigma_{t})_{t\geqslant 0}\in L^{1}(B_{t})\,|\,M_{t}=M_{0}+\int_{0}^{t}\sigma_{s}\,\mathrm{d}B_{s}\implies\right.$
$\hskip 145.10922ptM_{0}\sim\mu,M_{1}\sim\nu,(M_{t})\text{ is a martingale}\Bigg{\}},$ (1.1)
where $(B_{t})_{t\geqslant 0}$ is a standard Brownian motion and $(\mu,\nu)$ are convexly ordered probability measures with finite second moments. If $(\mu,\nu)$ is irreducible in the sense of [2], Appendix A.1, the solutions to the MBBP are called $(\mu,\nu)$-stretched Brownian motions, and have the form $\mathbb{E}[F(B_{1}+Y_{0})\,|\,\mathcal{F}_{t}]$, where $(\mathcal{F}_{t})$ is the filtration generated by $(B_{t}+Y_{0})$. The function $F:\mathbb{R}\rightarrow\mathbb{R}$ and the random variable $Y_{0}$ must be selected such that it is independent of $(B_{t})_{t>0}$, $F(B_{1}+Y_{0})\sim\nu$, and $\mathbb{E}[F(B_{1}+Y_{0})\,|\,Y_{0}]\sim\mu$. In the case where $(\mu,\nu)$ have irreducible components, the solution to the MBBP is a stretched Brownian motion on each irreducible component, see Theorem 3.1 in [1]. Stretched Brownian motions are interpolating martingales between the given measures $\mu$ and $\nu$. Another weak solution to the martingale interpolation problem is given by Schrödinger’s volatility models (SVMs) [10]. Inspired by the entropic relaxation of optimal transport ([18], [24]), also known as Schrödinger’s problem in specific cases [27], SVMs rely on a measure change. Consider two, possibly correlated, Brownian motions $(B_{t})_{t\geqslant 0}$ and $(\tilde{B}_{t})_{t\geqslant 0}$, under a measure $\mathbb{P}$, along with the system of SDEs on a real interval $[T_{0},T_{1}]$
$\,\mathrm{d}M_{t}=a_{t}\,M_{t}\,\mathrm{d}\widetilde{B}_{t},$
$\,\mathrm{d}a_{t}=b(a_{t})+c\left(a_{t}\right)\text{d}B_{t},$ (1.2)
where the real functions $b$ and $c$ are given, and are such that $(a_{t})_{t\in[T_{0},T_{1}]}$ is an Itô process. The goal is to change the measure $\mathbb{P}$ to an equivalent one, such that the process $(M_{t})_{0\leqslant t\leqslant 1}$ matches the given measures $\mu$ and $\nu$ at times zero and one, respectively, while being a true martingale. If several equivalent measures exist that satisfy these constraints, one must select the one that is closest to $\mathbb{P}$ in terms of the Kullback-Leibler divergence. By Girsanov’s theorem, changing $\mathbb{P}$ to an equivalent measure $\mathbb{Q}$ transforms the system of SDEs as follows:
$\,\mathrm{d}M_{t}=a_{t}M_{t}(\text{d}\widetilde{B}_{t}+\widetilde{\lambda}_{t}\,\mathrm{d}t),$
$\,\mathrm{d}a_{t}=b(a_{t})+c\left(a_{t}\right)(\text{d}B_{t}+\lambda_{t}\,\mathrm{d}t).$ (1.3)
One must then find the drifts $(\tilde{\lambda}_{t})$ and $(\lambda_{t})$, generated by the same measure change, that minimize the functional $\mathbb{E}_{\mathbb{Q}}[\int_{T_{0}}^{T_{1}}\widetilde{\lambda}^{2}_{s}+\lambda^{2}_{s}\,\mathrm{d}s]$ while satisfying the constraints $M_{0}\sim\mu$ and $M_{1}\sim\nu$, and so that $(M_{t})_{t\in[T_{0},T_{1}]}$ be a martingale.
@@
In this paper, inspired by the information-based approach, see [7] and [8], we construct a class of martingales that match, almost surely, any set of convexly ordered random variables, indexed by an arbitrary set of fixed times. The construction of such interpolating martingales, called filtered arcade martingales (FAMs), need two main ingredients: a discrete-time martingale (for example, a solution to an MOT problem) and an interpolating process that we call the randomized arcade process (RAP).
@@
In Section 2, we first develop the arcade processes (AP). Defined as a functional of a given stochastic process, called the driver, APs are sample-continuous stochastic processes that interpolate between zeros on the whole probability space, i.e., omega by omega. To achieve the interpolation, APs rely on deterministic functions, the interpolating coefficients. APs may be viewed as multi-period anticipative stochastic bridges. We study their properties and focus on Gaussian APs, which play an important role in the construction of FAMs. Starting from a Gauss-Markov driver, we show that it is always possible to construct a Markovian AP by utilizing the covariance structure of the driver. The resulting AP from this procedure is called a standard AP. This process is not unique, since one can construct infinitely many Markovian APs that are driven by the same Gauss-Markov process.
@@
Section 3 treats an additive randomization of APs, that is, an AP added to a stochastic process that interpolates deterministically between the given random variables. We term such processes randomized arcade processes (RAPs), which can be thought of as a sum of a signal function and a noise process. By construction, a RAP can match any random variables on the whole probability space at an arbitrary sequence of fixed times, i.e, it is a stochastic interpolator between target random variables in the strong sense. The notion of the Markov property does not suit RAPs in general, since their natural filtration contains the $\sigma$-algebras generated by each previously matched target random variables. For this reason, we introduce a weaker notion, the conditionally-Markov property, and go on to show under which conditions a RAP is conditionally-Markov.
@@
In the fourth section, we introduce the filtered arcade martingales (FAMs). A FAM is defined by the conditional expectation of the final target random variable, given the information generated by a RAP and, hence, inherits the filtering framework of the information-based approach that appears in [19], [7], [8]: the signal is the final target random variable and it can only be observed through a noisy version, the RAP. FAMs are tractable, thanks to the conditionally-Markov property enjoyed by the underlying RAPs, and can be simulated using Bayes formula. Applying Itô’s formula, we derive the stochastic differential equation satisfied by a FAM, which reveals the structure of the underlying innovations process that is adapted to the filtration generated by the respective RAP. The same ideas are applied to introduce the filtered arcade reverse-martingales (FARMs), constructed using reverse-filtrations of RAPs. Finally, as an application of the this theory, we briefly discuss a similar problem to martingale optimal transport, that incorporates noise in the optimization process, inspired by the entropic regularization of optimal transport and Schrödinger’s problem. This approach selects the martingale coupling that maximizes the expectation of the cumulative cost in time between a target random variable and its associated FAM, for a given RAP.
@@
In what follows, let $n\in\mathbb{N}_{0}$ and consider the collection of fixed dates $\{T_{i}\in\mathbb{R}\,|\,i=0,1,\ldots,n\}$ such that $0\leqslant T_{0}<T_{1}<T_{2}<\ldots<T_{n}<\infty$, and introduce the ordered sets
@@
Let $(\Omega,\mathcal{F},\mathbb{P})$ be a probability space, $\left(D_{t}\right)_{t\in[T_{0},T_{n}]}$ a sample-continuous stochastic process such that $\mathbb{P}[D_{t}=0]<1$ whenever $t\in(T_{0},T_{n})_{*}$, and $X$ an $\mathbb{R}^{n+1}$-valued random vector that is independent of $(D_{t})$.
@@ 2 Arcade processes
We construct stochastic processes on $[T_{0},T_{n}]$, as a functional of $(D_{t})$, which match exactly 00 (for all $\omega\in\Omega$) at the given times $\{T_{0},T_{n}\}_{*}$. The first step is to introduce deterministic functions called interpolating coefficients.
@@
An arcade process (AP), denoted $(A_{t}^{(n)})_{t\in[T_{0},T_{n}]}$, on the partition $\{T_{0},T_{n}\}_{*}$ is a stochastic process of the form
$A_{t}^{(n)}:=D_{t}-\sum_{i=0}^{n}f_{i}(t)D_{T_{i}},$ (2.1)
@@
We observe that $A_{T_{0}}^{(n)}=A_{T_{1}}^{(n)}=\ldots=A_{T_{n}}^{(n)}=0$ by construction, for all $\omega\in\Omega$. That is, $(A_{t}^{(n)})_{t\in[T_{0},T_{n}]}$ strongly interpolates between zeros at times $t\in\{T_{0},T_{n}\}_{*}$.
@@
For $n=1$, ${f_{0}(t)=\frac{t-T_{1}}{T_{0}-T_{1}}}$, ${f_{1}(t)=\frac{t-T_{0}}{T_{1}-T_{0}}}$, the AP driven by a standard Brownian motion $\left(B_{t}\right)_{t\geqslant 0}$ is the anticipative Brownian bridge on $[T_{0},T_{1}]$,
@@
Example 2.4.
For $n>1$, we can generalize the anticipative Brownian bridge by taking
$f_{0}(t)=\frac{T_{1}-t}{T_{1}-T_{0}}\mathds{1}_{[T_{0},T_{1}]}(t),\quad f_{n}(t)=\frac{t-T_{n-1}}{T_{n}-T_{n-1}}\mathds{1}_{(T_{n-1},T_{n}]}(t),$ (2.3)
and
$f_{i}(t)=\frac{t-T_{i-1}}{T_{i}-T_{i-1}}\mathds{1}_{(T_{i-1},T_{i}]}(t)+\frac{T_{i+1}-t}{T_{i+1}-T_{i}}\mathds{1}_{(T_{i},T_{i+1}]}(t),\quad\text{for }i=1,\ldots,n-1.$ (2.4)
We call this AP the stitched Brownian AP, for it can be written as
$A_{t}^{(n)}=\left\{\begin{array}[]{ll}B_{t}-\frac{T_{1}-t}{T_{1}-T_{0}}B_{T_{0}}-\frac{t-T_{0}}{T_{1}-T_{0}}B_{T_{1}},&\mbox{if }t\in[T_{0},T_{1}),\\ \\ B_{t}-\frac{T_{2}-t}{T_{2}-T_{1}}B_{T_{1}}-\frac{t-T_{1}}{T_{2}-T_{1}}B_{T_{2}},&\mbox{if }t\in[T_{1},T_{2}),\\ \,\vdots\\ B_{t}-\frac{T_{n}-t}{T_{n}-T_{n-1}}B_{T_{n-1}}-\frac{t-T_{n-1}}{T_{n}-T_{n-1}}B_{T_{n}},&\mbox{if }t\in[T_{n-1},T_{n}].\end{array}\right.$ (2.5)
Figure 1: Paths simulation of a stitched Brownian AP with $n=5$, using the equidistant partition $\{T_{i}=2i\,|\,i=0,1,\ldots,5\}$, c.f., Example 2.4.
Example 2.5.
Another way of generalizing the anticipative Brownian bridge to obtain an AP driven by Brownian motion is by using Lagrange’s polynomial interpolation, that is
$f_{i}(t)=\prod\limits_{k=0,k\neq i}^{n}\frac{T_{k}-t}{T_{k}-T_{i}}\quad\text{for }i=0,\ldots,n.$ (2.6)
$A_{t}^{(n)}=B_{t}-\sum_{i=0}^{n}\prod\limits_{k=0,k\neq i}^{n}\frac{T_{k}-t}{T_{k}-T_{i}}B_{T_{i}}.$ (2.7)
More generally, the Lagrange AP driven by a stochastic process $(D_{t})$ has the form
$A_{t}^{(n)}=D_{t}-\sum_{i=0}^{n}\prod\limits_{k=0,k\neq i}^{n}\frac{T_{k}-t}{T_{k}-T_{i}}D_{T_{i}}.$ (2.8)
Figure 2: Paths simulation of a Lagrange-Brownian AP with $n=5$, using the equidistant partition $\{T_{i}=2i\,|\,i=0,1,\ldots,5\}$, c.f., Example 2.5.
The Lagrange APs inherit Runge’s phenomenon from their interpolation coefficients. When $n$ is large, the AP oscillates around the edges of the interval.
Figure 3: Paths simulation of a Lagrange-Brownian AP with $n=10$, using the equidistant partition $\{T_{i}=i\,|\,i=0,1,\ldots,10\}$, illustrating Runge’s phenomenon, c.f., Example 2.5.
One can control this effect by applying a transformation to the interpolating coefficients. For instance, the map $x\mapsto|x|^{2(1-|x|)}$ applied to each of the interpolating coefficients
$f_{i}(t)=\prod\limits_{k=0,k\neq i}^{n}\frac{T_{k}-t}{T_{k}-T_{i}}$
for $i=0,\ldots,n$, yields another set of interpolating coefficients $\tilde{f}_{0},\ldots,\tilde{f}_{n}$ which do no longer suffer from Runge’s phenomenon.
Figure 4: Paths simulation of a Brownian AP with $n=10$ and interpolating coefficients $\tilde{f}_{0},\ldots,\tilde{f}_{10}$, using the equidistant partition $\{T_{i}=i\,|\,i=0,1,\ldots,10\}$, c.f., Example 2.5.
Example 2.6.
Elliptic APs have interpolating coefficients given by
$f_{0}(t)=\sqrt{1-\left(\frac{t-T_{0}}{T_{1}-T_{0}}\right)^{2}}\mathds{1}_{[T_{0},T_{1}]}(t),\quad f_{n}(t)=\sqrt{1-\left(\frac{t-T_{n}}{T_{n}-T_{n-1}}\right)^{2}}\mathds{1}_{(T_{n-1},T_{n}]}(t),$ (2.9)
$f_{i}(t)=\sqrt{1-\left(\frac{t-T_{i}}{T_{i}-T_{i-1}}\right)^{2}}\mathds{1}_{(T_{i-1},T_{i}]}(t)+\sqrt{1-\left(\frac{t-T_{i}}{T_{i+1}-T_{i}}\right)^{2}}\mathds{1}_{(T_{i},T_{i+1}]}(t),\quad\text{for }i=1,\ldots,n-1.$ (2.10)
Figure 5: Paths simulation of an elliptic-Brownian AP with $n=5$, using the equidistant partition $\{T_{i}=2i\,|\,i=0,1,\ldots,5\}$, c.f., Example 2.6.
The expectation and the covariance of an AP, when they exist, are fully characterised by the driver and the interpolating coefficients.
@@
$\mu_{A}(t):=\mathbb{E}[A_{t}^{(n)}]=\mu_{D}(t)-\sum\limits_{i=0}^{n}f_{i}(t)\mu_{D}(T_{i}),$ (2.11)
$\sigma_{A}^{2}(t):=\mathrm{Var}{[A_{t}^{(n)}]}=\sigma_{D}^{2}(t)+\sum\limits_{i=0}^{n}f_{i}^{2}(t)\sigma_{D}^{2}(T_{i})-2f_{i}(t)K_{D}(t,T_{i})$
@@
$K_{A}(s,t):=\mathrm{Cov}[{A_{s}^{(n)},A_{t}^{(n)}}]=K_{D}(s,t)-\sum\limits_{i=0}^{n}f_{i}(t)K_{D}(s,T_{i})+f_{i}(s)K_{D}(t,T_{i})$
@@
Markov APs will play a crucial role in the construction of filtered arcade martingales. A useful property of these processes is the following:
@@
In particular, if $(A_{t}^{(n)})$ is Markov and has a covariance function, then $\mathrm{Cov}(A_{s}^{(n)},A_{t}^{(n)})=0$ whenever $s\in[T_{i},T_{i+1}]$, $t\in[T_{j},T_{j+1}]$ and $j\neq i$.
If we think of APs as noise processes, it is natural to study their Gaussian subclass. But instead of looking at all APs that are also Gaussian processes, we focus on APs driven by Gaussian processes.
@@
An AP $(A_{t}^{(n)})_{t\in[T_{0},T_{n}]}$ is said to be a Gaussian AP if its driver $(D_{t})$ is a Gaussian stochastic process on $[T_{0},T_{n}]$.
Not all APs that are Gaussian processes are Gaussian APs, though all Gaussian APs are Gaussian processes. For example, take $D_{t}=B_{t}+tY$, where $(B_{t})$ is a Gaussian process and $Y$ a non-Gaussian random variable, and $A_{t}=D_{t}-(1-t)D_{0}-tD_{1}$. The driver is not Gaussian while the AP is a Gaussian process. Next, we give a first result on the Markov property of Gaussian APs.

### v3

@@ Abstract
Arcade processes are a class of continuous stochastic processes that interpolate in a strong sense, i.e., omega by omega, between zeros at fixed pre-specified times. Their additive randomisation allows one to match any finite sequence of target random variables, indexed by the given fixed dates, on the whole probability space. The randomised arcade processes (RAPs) can thus be interpreted as a generalisation of anticipative stochastic bridges. The filtrations generated by these processes are utilised to construct a class of martingales that interpolate between the given target random variables. These so-called filtered arcade martingales (FAMs) are almost-sure solutions to the martingale interpolation problem and reveal an underlying stochastic filtering structure. In the special case of conditionally Markov randomised arcade processes, the dynamics of FAMs are informed by Bayesian updating. The same ideas are applied to filtered arcade reverse-martingales, which are constructed in a similar fashion, using reverse-filtrations of RAPs, instead. Several explicit examples for RAPs and FAMs are provided and simulated. This paper concludes with an outlook on potential connections between FAMs and martingale optimal transport, and related applications.
keywords:
Arcade processes, stochastic interpolation, stochastic bridges, martingale interpolation, nonlinear stochastic filtering, information-based approach, Markov processes, Bayesian updating, martingale optimal transport.
[…]
@@ 1 Introduction
Strassen’s and Kellerer’s theorems do not provide a method for constructing the matching martingales. Many techniques inspired by Kellerer’s setting, where one looks to construct a martingale that matches a given peacock [12]: Skorokhod’s embedding problem, Brownian random time-changes, local volatility models, scaling peacock methods, etc., see [23]. Since the target is a dynamical measure, the matching is produced in a weak sense, and the martingale evolves in continuous time. On the other hand, the matching in Strassen’s setting can be weak or almost sure, depending on the nature of the targets (measures or random variables), and the martingale is in discrete time. A well-known approach for this task is martingale optimal transport (MOT), see, e.g., [2] and [3].
@@
Our interest lies at the intersection of Strassen’s and Kellerer’s settings: how to construct a continuous-time martingale that matches, in law or almost surely, a given finite set of convexly ordered random variables, at pre-specified dates? This is for instance of interest in a financial setting, where knowledge of the prices of vanilla call and put options provides an implied distribution for the underlying asset price at future fixed dates under a no-arbitrage requirement. This is an example of what is called the martingale interpolation problem, since one wishes to interpolate with a continuous-time martingale between the components of a discrete-time martingale. If one put the martingale condition aside, then bridging the gap between a coupling and an interpolating measure is usually a manageable proposition. Imposing the martingale property on the interpolating process is the challenging part of this problem.
@@
In this paper, inspired by the information-based approach, see [7] and [8], we construct a class of martingales that match, almost surely, any set of convexly ordered random variables, indexed by an arbitrary set of fixed times. The construction of such interpolating martingales, called filtered arcade martingales (FAMs), needs two main ingredients: a discrete-time martingale (for example, a solution to an MOT problem) and an interpolating process that we call the randomised arcade process (RAP).
Before we proceed with a section-by-section summary of the content of this paper, we next emphasize the mathematical and scientific contributions of the new approach for stochastic interpolation proposed here.
  1. 1.
    A new class of sample-continuous stochastic processes is constructed that interpolate strongly between zeros, that is, omega by omega, called arcade processes. In general, these are no Markov processes. However, sufficient conditions are established under which an AP is a Markov process.
  2. 2.
    Strong interpolation between a vector of random variables with arbitrary distribution is obtained by an additive extension of arcade processes—the randomised arcade processes. These processes feature a so-called signal process and an arcade process, the latter being the noise process. RAPs are no Markov processes. However, a special class of RAPs is identified for which a conditionally Markov property is introduced. Conditionally Markov randomised arcade processes is a class of processes that allows for closed-form expressions for the dynamics of the ensuing martingale interpolation.
  3. 3.
    Filtered arcade martingales, a class of processes which almost surely interpolate between the random variables (in convex order) linked by the randomised arcade processes, are constructed, whereby the FAMs are adapted to the filtration generated by the associated randomised arcade process. In the case that the randomised arcade process posses the conditional-Markov property, it is shown that the filtered arcade martingale can be expressed in closed form by an application of the Bayes formula. A similar construction method is produced to build filtered arcade reverse-martingales (FARMs), which are constructed in a similar fashion, using reverse-filtrations of RAPs.
  4. 4.
    We establish the connection between filtered arcade martingales and stochastic filtering problems, for which an explicit solution can be found in the case that the underlying randomised arcade process is conditionally Markov with respect to its natural filtration.
  5. 5.
    The almost-sure construction method of the interpolating filtered arcade martingales allows for a level of flexibility to search for the optimal martingale coupling of the random variables in-between which the filtered arcade martingale interpolates. This property of FAMs’ leads to the investigation of filtered arcade martingales in the context of martingale optimal transport which we discuss at the end of this paper in the outlook pointing to separate, ongoing research.
@@
Following this introduction, in Section 2, we begin with constructing the arcade processes (AP). Defined as a functional of a given stochastic process, called the driver, APs are sample-continuous stochastic processes that interpolate between zeros on the whole probability space, i.e., omega by omega. To achieve the interpolation, APs rely on deterministic functions, the interpolating coefficients. APs may be viewed as multi-period anticipative stochastic bridges. We study their properties and focus on Gaussian APs, which play an important role in the construction of FAMs. Starting from a Gauss-Markov driver, we show that it is always possible to construct a Markovian AP by utilising the covariance structure of the driver. The resulting AP from this procedure is called a standard AP. This process is not unique, since one can construct infinitely many Markovian APs that are driven by the same Gauss-Markov process.
@@
Section 3 treats an additive randomisation of APs, that is, an AP added to a stochastic process that interpolates deterministically between the given random variables. The randomised arcade processes (RAPs) can be thought of as a sum of a signal function and a noise process. By construction, a RAP can match any random variables on the whole probability space at an arbitrary sequence of fixed times, i.e, it is a stochastic interpolator between target random variables in the strong sense. The notion of the Markov property does not suit RAPs in general, since their natural filtration contains the $\sigma$-algebras generated by each previously matched target random variables. For this reason, we introduce a weaker notion, the conditionally-Markov property, and go on to show under which conditions a RAP is conditionally Markov.
@@
In the fourth section, we introduce the filtered arcade martingales (FAMs). A FAM is defined by the conditional expectation of the final target random variable, given the information generated by a RAP and, hence, inherits the filtering framework of the information-based approach that appears in [21], [7], [8]: the signal is the final target random variable and can only be observed through a noisy process, the RAP. FAMs are tractable, thanks to the conditionally-Markov property enjoyed by the underlying RAPs, and can be simulated using Bayes’ formula. Applying Itô’s formula, we derive the stochastic differential equation satisfied by a FAM, which reveals the structure of the underlying innovations process that is adapted to the filtration generated by the associated RAP. The same ideas are applied to introduce the filtered arcade reverse-martingales (FARMs), constructed using reverse-filtrations of RAPs. Finally, as an application of this theory, we briefly discuss a similar problem to martingale optimal transport that incorporates noise in the optimisation process, inspired by the entropic regularisation of optimal transport and Schrödinger’s problem. This approach selects the martingale coupling that maximises the expectation of the cumulative cost in time between a target random variable and its associated FAM, for a given RAP.
@@
Throughout this paper, unless specified otherwise, let $n\in\mathbb{N}_{0}$ and consider the collection of fixed dates $\{T_{i}\in\mathbb{R}\,|\,i=0,1,\ldots,n\}$ such that $0\leqslant T_{0}<T_{1}<T_{2}<\ldots<T_{n}<\infty$, and introduce the ordered sets
@@
Let $(\Omega,\mathcal{F},\mathbb{P})$ be a probability space, $\left(D_{t}\right)_{t\in[T_{0},T_{n}]}$ a sample-continuous stochastic process such that for all $c\in\mathbb{R}$, $\mathbb{P}[D_{t}=c]<1$ for all $t\in(T_{0},T_{n})_{*}$, and $X$ an $\mathbb{R}^{n+1}$-valued random vector that is independent of $(D_{t})$.
@@ 2 Arcade processes
We construct stochastic processes on $T_{0},T_{n}]$, as a functional of $(D_{t})$, which match exactly 0 (for all $\omega\in\Omega$) at the given times $\{T_{0},T_{n}\}_{*}$. The first step is to introduce deterministic functions called interpolating coefficients.
@@
An arcade process (AP) on the partition $\{T_{0},T_{n}\}_{*}$, denoted $(A_{t}^{(n)})_{t\in[T_{0},T_{n}]}$, is a stochastic process of the form
$A_{t}^{(n)}\mathrel{\mathop{\ordinarycolon}}=D_{t}-\sum_{i=0}^{n}f_{i}(t)D_{T_{i}},$ (2.1)
@@
We observe that $A_{T_{0}}^{(n)}=A_{T_{1}}^{(n)}=\ldots=A_{T_{n}}^{(n)}=0$ by construction for all $\omega\in\Omega$. That is, $(A_{t}^{(n)})_{t\in[T_{0},T_{n}]}$ strongly interpolates between zeros at times $t\in\{T_{0},T_{n}\}_{*}$.
@@
For $n=1$, $f_{0}(t)=\frac{t-T_{1}}{T_{0}-T_{1}}$}, $f_{1}(t)=\frac{t-T_{0}}{T_{1}-T_{0}}$}, the AP driven by a standard Brownian motion $\left(B_{t}\right)_{t\geqslant 0}$ is the anticipative Brownian bridge on $T_{0},T_{1}]$,
@@
Two generalisations of the Brownian bridge $(A_{t}^{(1)})_{t\in[T_{0},T_{1}]}$, shown in Example 2.3, to an $n$-arc arcade process $(A_{t}^{(n)})_{t\in[T_{0},T_{n}]}$ are given in A, where their properties are investigated and other examples are provided. We call the generalisation
$A_{t}^{(n)}=D_{t}-\sum_{i=0}^{n}\prod\limits_{k=0,k\neq i}^{n}\frac{T_{k}-t}{T_{k}-T_{i}}D_{T_{i}},$ (2.3)
driven by a stochastic process $(D_{t})$, the Lagrange AP due to the form of the interpolating coefficients.
One might wonder why we choose the timeline $T_{0},T_{1},\ldots,T_{n}$ and the interpolating coefficients $f_{0},f_{1},\ldots,f_{n}$ to be deterministic rather than stochastic. While the definition of APs is easily extended to a stochastic timeline and stochastic interpolating coefficients, APs are fundamentally designed as a building bloc for the martingales introduced in Section 5. Inspired by several applications, these martingales interpolate between random variables at fixed times. APs play a key role in their construction, which relies entirely on a deterministic timeline and deterministic interpolating coefficients. Furthermore, we want APs to be understood as a natural generalisation of anticipative bridges, which also use deterministic times and interpolating coefficients.
The expectation and covariance of an AP, when they exist, are fully characterised by the driver and the interpolating coefficients.
@@
$\mu_{A}(t)\mathrel{\mathop{\ordinarycolon}}=\mathbb{E}[A_{t}^{(n)}]=\mu_{D}(t)-\sum\limits_{i=0}^{n}f_{i}(t)\mu_{D}(T_{i}),$ (2.4)
$\sigma_{A}^{2}(t)\mathrel{\mathop{\ordinarycolon}}=\mathrm{Var}{[A_{t}^{(n)}]}=\sigma_{D}^{2}(t)+\sum\limits_{i=0}^{n}f_{i}^{2}(t)\sigma_{D}^{2}(T_{i})-2f_{i}(t)K_{D}(t,T_{i})$
@@
$K_{A}(s,t)\mathrel{\mathop{\ordinarycolon}}=\mathrm{Cov}[{A_{s}^{(n)},A_{t}^{(n)}}]=K_{D}(s,t)-\sum\limits_{i=0}^{n}f_{i}(t)K_{D}(s,T_{i})+f_{i}(s)K_{D}(t,T_{i})$
@@
Markov APs will play a crucial role in the construction of filtered arcade martingales. A useful property of these processes is the following.
@@
Proof.
Since $(A_{t}^{(n)})$ is 0 on $\{T_{0},T_{n}\}_{*}$, the proof follows immediately from the Markov property. ∎
In particular, if $(A_{t}^{(n)})$ is Markov, and it has a covariance function, then $\mathrm{Cov}(A_{s}^{(n)},A_{t}^{(n)})=0$ whenever $s\in[T_{i},T_{i+1}]$, $t\in[T_{j},T_{j+1}]$ and $j\neq i$.
If APs are considered being noise processes, it is natural to study their Gaussian subclass. Therefore, we introduce terminology that avoids confusion when referring to the driver of an AP that is Gaussian, since the driver of an AP is not uniquely determined. Consider the process $Y_{t}=B_{t}+tY$, where $(B_{t})$ is a Gaussian process, $Y$ a non-Gaussian random variable, and the AP $(A_{t}^{(1)})_{t\in[0,1]}$ on $\{0,1\}$ is given by
$A_{t}^{(1)}$ $=Y_{t}-(1-t)Y_{0}-tY_{1}$
$=B_{t}+tY-(1-t)B_{0}-t(B_{1}+Y)$
$=B_{t}-(1-t)B_{0}-tB_{1}.$ (2.11)
Both $(Y_{t})$ and $(B_{t})$ can be considered the driver of $(A_{t}^{(1)})$, which is a Gaussian process, but only one driver is Gaussian.
@@
An AP $(A_{t}^{(n)})_{t\in[T_{0},T_{n}]}$ is said to be a Gaussian AP if its driver $(D_{t})$ is a Gaussian stochastic process on $T_{0},T_{n}]$.
Under this definition, we know that $(Y_{t})$ cannot be referred to as the driver of the Gaussian AP $(A_{t}^{(1)})$ defined above, since it is not Gaussian. Next, we give a first result on the Markov property of Gaussian APs.